### v2

# A Matter of Annotation: An Empirical Study on In Situ and Self-Recall Activity Annotations from Wearable Sensors

Abstract
Research into the detection of human activities from wearable sensors is a highly active field, benefiting numerous applications, from ambulatory monitoring of healthcare patients via fitness coaching to streamlining manual work processes. We present an empirical study that compares 4 different commonly used annotation methods utilized in user studies that focus on in-the-wild data. These methods can be grouped into user-driven, in situ annotations - which are performed before or during the activity is recorded - and recall methods - where participants annotate their data in hindsight at the end of the day. Our study illustrates that different labeling methodologies directly impact the annotations’ quality, as well as the capabilities of a deep learning classifier trained with the data. We noticed that in situ methods produce less but more precise labels than recall methods. Furthermore, we combined an activity diary with a visualization tool that enables the participant to inspect and label their activity data. Due to the introduction of such a tool were able to decrease missing annotations and increase the annotation consistency, and therefore the F1-Score of the deep learning model by up to 8% (ranging between 82.1 and 90.4 % F1-Score). Furthermore, we discuss the advantages and disadvantages of the methods compared in our study, the biases they could introduce, and the consequences of their usage on human activity recognition studies as well as possible solutions.
Keywords Human-centered computing ubiquitous computing user-driven study annotations
Sensor-based activity recognition is one of the research fields of Pervasive Computing developed with enormous speed and success by industry and science and influencing medicine, sports, industry, and therefore the daily lives of many people. However, current available smart devices are mostly capable of detecting periodic activities like simple locomotions. To recognize more complex activities a multimodal sensor input, such as [1], and more complex recognition models are needed. Many of the published datasets are made in controlled laboratory environments. Such data does not have the same characteristics and patterns as data recorded in-the-wild. Data that belongs to similar classes but is recorded in an uncontrolled versus controlled environment can differ significantly since it contains more contextual information [2]. Furthermore, study participants tend to control their movements more while being monitored [3]. The recording of long-term and real-world data is a tedious, time-consuming, and therefore a non-trivial task. Researchers have various motivations to record such datasets but the technical hurdles are still high and problems during the annotation process occur regularly. With regards to Human Activity Recognition (HAR), recording a long-term dataset always presents the researcher with the problem of developing a methodology that, on one hand, allows precise labels and, on the other hand, does not unnecessarily burden or disturb the study participants. Relying only on self-recall methods, like writing an activity diary, e.g. [4], can result in imprecise time indications that do not necessarily correspond to the actual periods of an activity. Such incorrectly or noisy labeled data [5] leads to a trained model that is less capable of detecting activities reliably, due to unwanted temporal dependencies learned by wrongly annotated patterns [6].
We can see an emerging spotlight on real-world and long-term activity recognition and think that it will be one of the main research challenges that need to be put more in focus to overcome current limitations and be capable of recognizing complex day-to-day activities.
Such datasets, rely heavily on self-recall methods or using additional apps to track movements and set labels either automatically [7] or with the manual selection of a label [8]. Due to these hurdles, many researchers prefer to work with datasets from controlled over uncontrolled environments. As a consequence, only a very limited number of in-the-wild datasets have been published until now.
Contribution: Our study focuses on the evaluation of 4 different annotation methods for labeling data in-the-wild: In situ (lat. on site or in position) with a button on a smartwatch, in situ with the app Strava 111https://www.strava.com/ (an app that is available for iOS and Android smartphones), pure self-recall (writing an activity diary at the end of the day), and time-series assisted self-recall with the MAD-GUI [9], which displays the sensor data visually and allows to annotate it interactively. Our study was conducted with 11 participants, 10 males, and 1 female, over 2 weeks. Participants wore a Bangle.js Version 1222https://www.espruino.com/banglejs smartwatch on their preferred hand, used Strava, and completed self-recall annotations every evening.
In the first week, the participants were asked to write an activity diary at the end of the day without any helping material and additionally using two user-initiated methods (in situ button and in situ app) to manually set labels at the start and beginning of each activity. In the second week, the participants were given an additional visualization of the sensor data with an adapted version of the MAD-GUI annotation tool. With the help of this, participants then were instructed to label their data in hindsight with the activity diary as a mnemonic aid. Given labels from both weeks were compared to each other regarding the quality through visual inspection and statistical analysis with regard to the consistency and quantity of missing annotations across labeling methods.
Furthermore, we used a Shallow-DeepConv(LSTM) architecture, see Bock et al. [10] and Ordoñez et al. [11], and trained models with a leave-one-day-out cross-validation method of 6 previously selected subjects and each annotation method.
Impact: Annotating data, especially in real-world environments, is still very difficult and tedious. Labeling such data is always a trade-off between accuracy and workload for the study participants or annotators. We raise awareness among researchers to put more effort into exploring new annotation methods to overcome this issue. Our study shows that different labeling methodologies have a direct impact on the quality of annotations. With the deep learning analysis, we prove that this impacts the model capabilities directly. Therefore, we consider the evaluation of frequently used annotation methods for real-world and long-term studies to be crucial to give decision-makers of future studies a better base on which they can choose the annotation methodology for their study in a targeted way.
1 Materials and Methods
Our study is conducted with 11 participants, from which 10 are male and 1 is female. The participants are between 25 and 45 years old. Out of 11 participants, 6 are researchers in the field of signal processing and are used to read and work with sensor data. Participants were selected among acquaintances and colleagues. The study was conducted over 2 weeks while participants wore an open-source smartwatch on their wrist of choice. During the two-week study, the participants were instructed to use 4 different labeling methods in parallel, see Figure 1. In the first week they were asked to use the in situ button, in situ app, and pure self-recall methods. At the beginning of the 2nd week, we expanded the number of annotation methods with the time-series recall. This annotation method combines the activity diary with a graphical visualization of the participants’ daily data.
The Bangle.js smartwatch has 3 mechanical buttons on the right side of the case. These buttons are programmed to record the number of consecutive button presses per minute. The total number of button presses is stored with the given timestamp and can therefore be used to mark the beginning and end of an activity in the time-series.
In addition, the participants were asked to track their activities with the smartphone app Strava. Strava is an activity tracker that is available for Android and iOS and freely downloadable from the app stores. The user can choose from a variety of predefined labels and start recording. Recording an activity starts a timer that runs until the user stops it. The time as well as the GPS position of the user during the activity is tracked and saved locally.
The pure self-recall methods consist of writing an activity diary on a daily basis at the end of the day. The participants were explicitly told that they should only write down the activities that they still remember 2 hours after the measurement stopped.
The time-series recall method can be seen as a combination of an activity diary and a graphical representation of the raw sensor data. For visualization and labeling purposes, we provided the participants with an adapted version of the MAD-GUI. The GUI was published by Ollenschläger et al. [9] in 2022 and is a generic open-source Python package. Therefore, it can be integrated into one’s project. Our adaptions to the package are available for download from a GitHub repository333https://github.com/ahoelzemann/mad-gui-adaptions/. It contains changes to the data loader, the definition of available labels, and color settings for displaying the 3D raw data.
Annotation Guidelines: The participants were briefed to note daily returning activities (sports or activities of daily living) that are performed longer than 10 minutes roughly 2 hours after the recording stops. The name of the activity was chosen individually by the participant.
Participants decided individually at what time of the day the recording would start on the next day.
Each of these annotation methods represents a layer of annotation that is used for the visual, statistical, and deep learning evaluation. Figure 1 illustrates the overall concept.
2 Statistical Analysis
The labels were statistically analyzed based on their consistency using the Cohen score as well as the number of missing annotations across all methods. The Cohen score describes the agreement between two annotation methods, which is defined as follows (see [12] and [13]). Where is the observed agreement ratio and is the expected agreement if both annotators assign labels randomly. The score shows how uniform two different annotators labeled the same data. For calculation purposes, an implementation provided by Scikit-Learn [14], was used. Furthermore, missing annotations across methods are measured as the percentage of missing or incomplete annotations. The annotations of all methods were first compared with each other and matched based on the given time indications. Annotations that could not be assigned or were missing were marked accordingly and are the base for calculating this indicator, Figure 4 visualizes this. We used a similar representation as [15] to visualize the matches among labeling methods. In this study, the authors compared genome annotations labeled by different annotators with regard to their error scores between different annotators.
3 Effects on Deep Learning Performance
The deep learning analyses are performed using the DeepConvLSTM architecture [11] which is based on a Keras implementation of [16]. We did not perform hyperparameter tuning because it would involve a considerable amount of additional workload, since we trained 64 models independently during the evaluation. We therefore decided to opt out of the architecture with regards to efficiency rather than optimal classification results. Additionally, we don’t expect that the actual experiment - evaluating different annotation methods - would benefit from hyperparameter tuning or gain any significant information and insights. Instead, we use the default hyperparameters provided by the authors. These are depicted in the Figure 2. Furthermore, we reduce the number of LSTM layers to one and instead increase the number of hidden units of the only LSTM layer to 512. According to [10], this modification decreases the runtime up to 48 % compared to a two-layered DeepConvLSTM while significantly increasing the overall classification performance on 4/5 publicly available datasets: [17], [18], [1], [19], [20]. LSTM-Layers in general are important if the dataset contains sporadic activities [6]. However, our dataset does not and our evaluation aims to identify long periods of periodic activities, like walking or running. For this reason, we can conclude that additional LSTM layers are not needed.
The implementation of [16] incorporates BatchNormalization layers after each Convolutional layer, as well as MaxPooling for the transition between the final convolutional block and the LSTM layer, and a Dropout layer before classification. Each Convolutional layer employs a ReLU activation function. The inclusion of the BatchNormalization layers serves to accelerate training and mitigate the detrimental effects of internal covariate shift, as discussed further in [21]. Before training the neural network we apply two preprocessing steps to the data. These are an upsampling to a constant 25 Hz due to small deviations in the device sampling rate as well as a rescaling of the accelerometer data between -1 and 1.
Leave-One-Day-Out Cross-Validation: Figure 3 shows the train- and test-setting for the deep learning model. For every study participant and every week, a personalized model has been trained and tested.
We adapted the leave-one-subject-out strategy to fit our needs. Instead of using one participant for testing, we used one day of the week for testing and trained on every other day. The dataset consists of a major void-class and a small number of samples per activity class and participant. To counteract an above-average large void-class, we trained our model with balanced class weights. Not limiting the participants in their choice of daily performed activities as well as well as not specifying predefined activity labels resulted in very unique sets of activities per study participant. Due to these circumstances, we cannot expect that it is possible to train a model that is capable of generalizing across participants and days. Every participant comes with their specific patterns to perform an activity. Furthermore, due to the in-the-wild recording setup, the intra-class differences [22] for comparatively simple activities, walking or running can be significant. The impact of different labeling methods is therefore expected to be more present, and hence more visible, in a personalized than it would in a generalized model.
Postprocessing & Classification: We classified the data based on a sliding window with a length of 2 seconds (50 samples). However, we are looking for longer periods of reoccurring activities and decided therefore to apply a jumping window of 5 minutes that applies a majority vote to the given time period. The activity with the most instances in those 5 minutes is set for the whole window.
4 Results
Our participants were asked to annotate activities carried out in their daily lives performed for more than 10 min. We didn’t limit the participants to a predefined set of classes. They decided independently which labels they would like to set for certain periods. After normalizing the labels, e.g. changing “going for a walk” to “walking” etc., the set of labels as given by the participants contains the following 23 (22 + void) different labels: laying, sitting, walking, running, cycling, bus_driving, car_driving, vacuum_cleaning, laundry, cooking, eating, shopping, showering, yoga, sport, playing_games, desk_work, guitar_playing, gardening, table_tennis, badminton, horse_riding. Every sample that wasn’t specifically labeled is classified as void.
4.1 Missing Annotations and Consistency Across Methods
Missing Annotations and the consistency of labels set over one week varied greatly depending on the study participant. However, tendencies concerning specific methods are observable.
Method , pressing the situ button on the smartwatch’s case, was not consistently used by every participant. Furthermore, this method carries the risk that either setting one of the two markers (start or end) is forgotten. An annotation where one marker is missing becomes therefore obsolete.
The app-assisted annotation method , for which we used the app Strava, is well accepted among the participants who agreed with using third-party software. However, 4 participants, namely 74e4, 90a4, d8f2 and f30d did not use the app continuously or refused to use it completely due to concerns regarding their private data. Strava is a commercial app, that is freely available for download on the app stores, but it collects certain users’ metadata. To label a time period with Strava, the participant needs to (1) take the smartphone, (2) open the app, (3) start a timer, set a label, and (4) end the timer. This procedure contains significantly more steps than other methods. Therefore, the average value of missing annotations results in 46.40% (week 1) and 56.79% (week 2). One participant found the annotation process in general very tedious and therefore dropped out of the study. These data have been excluded from the dataset and the evaluation. Method pure self-recall, writing an activity diary, got well accepted by every participant. As Figure 4 shows and the results in Table 1 proof, it is overall the most complete annotation method with an average amount of missing annotations of 4.30 % for the first and 8.14 % for the second week.
Week 1 Subject 2b88 36fd 74e4 90a4 834b 4531 a506 d8f2 eed7 f30d fc25 Avg. in situ button 40 70.59 79.41 52.18 36.37 50 26.32 96.15 45.46 0 26.81 40.95 in situ app 13.30 5.89 97.06 100 5.00 0 36.84 92.30 22.73 100 4.35 43.40 pure self-recall 6.67 0 0 0 4.55 0 31.58 0 4.55 0 0 4.30 Week 2 in situ button 23.08 73.33 92.00 82.14 8.33 76.47 0 95.33 61.11 0 27.78 49.05 in situ app 61.54 6.67 100 89.29 8.33 35.30 100 79.16 33.33 100 11.11 56.79 pure self-recall 0 0 8.57 17.88 4.17 35.30 0 12.50 5.56 0 5.56 8.14 time-series recall 0 0 22.88 39.29 0 0 0 16.67 0 0 5.56 7.67
By introducing the MAD-GUI, participants were able to inspect their daily data, get insights into what patterns of specific classes look like, and label them interactively. With an average amount of missing annotations of 7.67%, this method became the most complete during the second study week.
Week, Day Direction 2b88 36fd 4531 74e4 834b 90a4 a506 d8f2 eed7 f30d fc25 Avg. 1, 1 C/W C/W 0.32 0.69 0.0 0.85 0.0 0.69 0.0 0.0 0.69 0.76 0.35 0.0 0.79 0.0 0.22 0.0 0.0 0.90 0.23 0.0 0.58 0.49 1, 2 C/W C/W 0.64 0.64 0.69 0.68 0.0 0.84 0.09 0.05 0.85 0.86 0.05 0.0 0.51 0.50 0.0 0.0 0.55 0.93 0.47 0.0 0.74 0.73 1, 3 C/W C/W 0.0 0.86 0.62 0.0 -0.03 -0.03 0.0 0.0 0.39 0.38 0.56 0.0 0.53 0.44 0.0 0.0 0.05 0.28 0.53 0.0 0.51 0.54 1, 4 C/W C/W 0.38 0.99 0.30 0.69 0.91 0.90 0.08 0.0 0.63 0.74 0.03 0.0 0.80 0.57 0.0 0.69 0.66 0.80 0.80 0.0 0.0 0.0 1, 5 C/W C/W 0.33 0.32 0.33 0.83 0.0 0.0 0.04 0.0 0.39 0.37 0.32 0.0 0.93 0.96 0.0 0.0 -0.03 -0.31 0.93 0.0 0.87 0.89 1, 6 C/W C/W 0.0 -0.14 0.0 0.96 0.75 0.71 0.07 0.0 0.0 0.15 0.34 0.0 0.67 -0.07 0.0 0.41 0.42 0.52 0.99 0.0 0.84 0.84 1, 7 C/W C/W 0.30 0.78 0.0 0.15 0.56 0.69 0.04 0.0 0.0 0.10 0.525 0.0 0.99 0.43 0.0 0.0 0.29 0.42 0.90 0.0 0.49 0.77 2, 1 C/W C/W C/W C/W C/W 0.30 0.45 0.85 0.39 0.53 0.56 0.77 0.76 0.43 0.61 0.36 0.37 0.56 0.48 0.45 0.10 0.0 0.10 0.43 0.0 0.51 0.57 0.48 0.74 0.70 0.0 -0.02 0.11 0.0 0.18 0.88 0.0 0.90 0.98 0.0 0.0 0.63 0.78 0.0 0.71 0.41 0.51 0.74 0.58 0.57 0.89 0.0 0.86 0.82 0.0 0.85 0.81 0.46 0.58 0.55 2, 2 C/W C/W C/W C/W C/W 0.82 0.84 -0.02 -0.02 -0.02 0.21 0.75 0.82 0.11 0.83 0.91 0.93 0.62 0.66 0.63 0.05 0.0 0.09 0.38 0.0 0.47 0.90 0.84 0.47 0.92 0.03 0.0 0.09 0.77 0.0 0.70 0.0 0.70 1.0 0.0 0.0 0.87 0.83 0.0 0.96 0.29 0.59 0.56 0.43 0.71 0.74 0.0 0.85 0.70 0.0 0.36 0.81 0.46 0.45 0.46 2, 3 C/W C/W C/W C/W C/W 0.70 0.66 0.68 0.91 0.74 0.44 0.44 0.54 0.53 0.53 0.0 0.98 0.62 0.0 0.82 0.0 0.0 0.91 0.0 0.0 0.62 0.72 0.49 0.77 0.74 0.0 0.0 -0.18 0.0 0.0 0.90 0.0 0.90 1.0 0.0 0.0 0.47 0.86 0.0 0.60 0.28 0.36 0.39 0.88 0.79 0.99 0.0 0.98 0.96 0.0 0.68 0.82 0.58 0.41 0.65 2, 4 C/W C/W C/W C/W C/W 0.45 0.14 0.17 0.71 0.82 0.0 0.85 0.86 0.0 0.51 0.83 0.83 0.84 0.86 0.86 0.0 0.0 0.90 0.0 0.0 -0.02 -0.02 -0.02 0.80 0.80 0.05 0.0 0.0 0.47 0.0 0.64 0.0 0.64 1.0 0.0 0.0 0.0 0.84 0.0 0.0 0.23 0.26 0.38 0.68 0.50 0.67 0.0 0.86 0.66 0.0 0.78 0.87 0.92 0.76 0.85 2, 5 C/W C/W C/W C/W C/W 0.59 0.0 0.48 0.5 0.0 0.0 0.41 0.47 0.0 0.34 0.0 0.77 0.76 0.0 0.89 0.0 0.0 0.16 0.0 0.0 0.28 0.28 0.26 0.82 0.83 0.09 0.01 0.09 0.94 0.60 0.60 0.0 0.59 0.99 0.0 0.0 0.54 0.82 0.0 0.49 0.0 0.54 0.40 0.0 0.70 0.73 0.0 0.43 0.38 0.0 0.95 0.92 0.47 0.46 0.44 2, 6 C/W C/W C/W C/W C/W 0.48 0.0 0.47 0.98 0.0 0.0 0.85 0.86 0.0 0.88 0.90 0.96 0.92 0.95 0.89 0.0 0.0 0.77 0.0 0.0 0.39 0.39 0.30 0.47 0.62 0.0 0.02 0.0 0.0 0.69 0.73 0.0 0.72 0.99 0.0 0.0 0.55 0.83 0.0 0.43 0.0 0.83 0.86 0.0 0.95 0.20 0.0 0.74 0.46 0.0 0.86 0.87 0.86 0.83 0.82 2, 7 C/W C/W C/W C/W C/W 0.0 0.0 0.96 0.0 0.0 0.0 0.41 0.47 0.0 0.80 0.0 0.76 0.72 0.0 0.93 0.0 0.0 0.0 0.0 0.0 0.30 0.30 0.24 0.79 0.78 0.0 0.0 -0.01 0.0 0.0 0.73 0.0 0.42 0.67 0.0 0.40 0.0 0.79 0.46 0.0 0.43 0.14 0.33 0.71 0.46 0.91 0.0 0.92 0.94 0.0 0.86 0.86 0.84 0.78 0.78 Avg. Week 2 C/W C/W C/W C/W C/W 0.48 0.30 0.51 0.50 0.30 0.17 0.64 0.68 0.15 0.64 0.43 0.80 0.72 0.42 0.78 0.02 0.0 0.12 0.42 0.0 0.36 0.45 0.37 0.69 0.78 0.02 0.0 0.01 0.31 0.21 0.74 0.0 0.70 0.94 0.0 0.0 0.44 0.82 0.07 0.46 0.23 0.46 0.33 0.47 0.67 0.73 0.0 0.81 0.70 0.0 0.76 0.85 0.66 0.61 0.65 0.39 0.36 0.52 0.48 0.41
Table 2 shows the resulting Cohen scores. Due to the constraint that only one labeling method can be compared to a second one and since, according to Table 1, the most consistent annotation methods are pure self-recall and time-series recall, we used these methods as our baseline and compared them with every other method used in the study. The second column indicates the comparison direction. The abbreviations used in this column are defined as follows: ( C/W ) pure self-recall compared with in situ button, ( C/W ) pure self-recall compared with in situ app and ( C/W ) pure self-recall compared with time-series recall. The direction ( C/W ) is not explicitly included since Cohens is bidirectional and both directions result in the same score. The score indicates how similar two annotators, or in our study labeling methods, are to each other. The resulting score is a decimal value between -1.0 and 1.0, where -1.0 means that the two annotators differ at most and 1.0 means complete similarity. 0.0 denotes that the target method was not used on that specific day.
Comparing the pure self-recall method with the in situ button and in situ app method we can see that the final results for weeks 1 and 2 are proximate to one another. Pure self-recall compared with the time-series recall results in the highest similarity of 0.52. The comparison between the time-series recall and the in situ button as well as the in situ app assisted annotations result in higher similarity than the prior comparison of pure self-recall vs. both methods and . This means that subjects rather agree to the timestamps of the in situ methods than to a self-written activity diary as soon as they can visually inspect the accelerometer data.
4.2 Visual Time-Series Analysis
Figure 5 shows exemplary the time-series of the sixth day of every participant’s second week.
The four bars that are visible above the accelerometer data are the labels set by the participants for every layer. The order is from bottom to top: in situ button, in situ app. pure self-recall, and time-series recall. Examples of labels that differ regarding the applied labeling method are marked with red boxes. The x-axis of every subplot represents roughly 8-9 hours of data. Most of the day was not labeled and is therefore categorized as void. However, such long periods often contain shorter periods of other activities, like walking. This makes it difficult to define a distinguishable void-class, which results in false positive classifications of non-void samples. Figure 5 visually shows that each participant labels his or her data very subjectively. The long green-labeled periods of participant 74e4 represent the class desk_work. The only other participant that used this label is 90a4. Since each of the study participants works in an office environment and thus conclusively works at a desk, we can assume that the same class is classified as void for all other study participants. This intra-class and inter-participant discrepancy becomes a problem whenever a model is trained that is supposed to generalize across individuals. To reduce these side effects and focus on the experiment itself, we decided to evaluate personalized models that take weekly data from participants into account.
The in situ button annotation is empty for 5 participants: eed7, 36fd, 74e4, 90a4 and d8f2. Labels are only partially set or missing entirely for this annotation method and we therefore assume that participants tend to forget to press the button on the smartwatch. Both tables, 1 and 2, support this assumption, as this labeling method shows a high percentage of missing annotations as well as a low Cohen score of 0.36% (week 1) and 0.39% (week 2). The pure self-recall method , visible on the 2nd upper layer, is often misaligned compared to the in situ methods as well as the time-series recall method . Participants tend to round up or down the start- and stop-time in steps of 5 or 10 minutes. For example, the annotations in Figure 5 given by the subjects 2b88, 834b or f30d, show such incorrectly annotated data. The pink color represents the class walking. With a closer look at the corresponding time-series data, one can see that the in situ button annotation (bottom layer) and time-series recall annotation (top layer) belongs to the typical periodic pattern of walking than the period labeled by pure self-recall.
A consistent reliable performance in all labeling methods can only be observed at the participants 4531 and fc25. Other participants like eed7, 36fd, 74e4 or a506 are very precise in their annotations across methods, but are missing at least one layer of labels. The complete collection of visualizations is available in our dataset repository444https://doi.org/10.5281/zenodo.7654684
4.3 Effects on Classification
The results of our deep learning evaluation555Detailed results for every participant included in our deep learning evaluation can be accessed online on the Weights & Biases platform: https://tinyurl.com/4vxvfaed. suggest that the annotation method chosen can have a crucial impact on the classification ability of a trained deep learning model. Depending on the chosen methodology, the average F1-Score results differ by up to 8%, as depicted in Figure 6.
In the first week, the in situ methodologies, button and app , generally perform better than the pure self-recall diary .
Week 1 Subject 2b88 a506 eed7 fc25 4531 834b Average in situ button 0,91 0,92 0,89 0,89 0,91 0,91 90,4 in situ app 0,92 0,60 0,91 0,84 0,93 0,92 85,5 pure self-recall 0,78 0,76 0,83 0,86 0,86 0,89 83,0 Week 2 in situ button 0,88 0,92 0,90 0,92 0,86 0,72 86,8 in situ app 0,91 na 0,92 0,94 0,85 0,88 90,1 pure self-recall 0,81 0,86 0,82 0,94 0,83 0,67 82,1 time-series recall 0,90 0,91 0,95 0,86 0,86 0,86 89,0
Study participants mostly correctly estimated the duration of an activity, but tended to round up or down the start and end times. The in situ methods are up to 8% better than the pure self-recall, although the amount of annotated data available, due to missing annotations, is significantly lower than for other methods. Although, we work with a dataset recorded in-the-wild, the deep learning results generally show a high F1-Score. This is untypical for such datasets but can be explained by the fact that the majority of the daily data are assigned to the void class. This leaves proportionally only a few samples that are crucial for determining the classification performance.
Even though the number of available annotations that have been labeled by the study participants using the time-series recall method is significantly higher with 92.33%, the average F1-Score is 1.1% lower (89.00%) than the results reached with the App Assisted method (90.1%). To understand this result it is crucial to look at Table 3 in detail and take meta-information about the participants into account. The participants mostly used their diary as a mnemonic aid for the graphical annotation method and tried to identify the corresponding periods in the acceleration data. The results of subjects 2b88, a506 and eed7 show that the performance of the classifier could be increased with graphical assistance. However, the F1-Score of 2b88 is 0.01% below the F1-Score of the in situ app assisted annotation method . These subjects have in common that they are already trained in interpreting acceleration data due to their prior knowledge and thus assign samples to specific classes more precisely.
Subjects fc25, 4531, and 834b, on the other hand, do not have prior knowledge. Apart from subject 834b, the deep learning results show that presenting a visualization to an untrained participant rather harmed than helped the classifier. If one looks at the visualizations of day 1 & 6, week 2 of fc25, see Figure 5 and 7, the labels set by the subject with the help of the graphical interface, it is comprehensible that this study participant tended to be rather confused by the graphical representation and therefore labeled the data incorrectly.
5 Discussion
In our 2-week long-term study, we recorded the acceleration data of 11 participants using a smartwatch and analyzed it visually, statistically, and using deep learning. The findings of the visual and statistical analysis were confirmed by the deep learning result. They show that the underlying annotation procedure is crucial for the quality of the annotations and the success of the deep learning model.
The in situ button method offers accuracy but brings the risk that the setting of a label is forgotten entirely or incompletely set. However, this method can be combined with additional on-device feedback or a smartphone app, so that greater accuracy and consistency of the annotation can be achieved. This involves a considerable implementation effort, which many scientists avoid because such projects, although of their significant value to the community, attract little attention in the scientific world. The use of existing, but often commercial, software and hardware is all too often accompanied by a loss of privacy. As our research has shown in passing, many users therefore shy away from using such products.
Through our investigation of the consistency of annotations between methodologies, we were able to show that participants in our study seem to prefer to write an activity diary (pure self-recall method ). This finding corresponds to what Vaizmann et al. [23] already points out. However, this method has the disadvantage that it can be imprecise, which is evident in the visualization of the data and annotations. Similarly, the activity diary methodology performed the least reliably among all methodologies, which has been confirmed by the deep learning model. Since the deep learning results using the in situ app annotations are almost similar to the results given by the time-series recall , even though the number of labeled samples is lower, it raises the question if a smaller set of high-quality annotations is more valuable for a classifier than a larger set of annotated data that comes with imprecise labels. This could mean that in future works we can reduce the amount of necessary training samples drastically if a certain annotation quality can be assured. However, this needs to be confirmed by further investigations.
Some participants reported that they found the support provided by the visual representation of the data helpful. The resulting Cohen scores strengthen this impression since the F1-Scores are much higher when we compare the time-series recall with both in situ methods vs. the pure self-recall. This indicates that as soon as the participants received a visual inspection tool, they tended to annotate data at similar time periods as through the in situ methods since they can easily identify periods of activity that roughly correspond to the execution time they remember. Our participants reported similar preferences, which led us to the conclusion that a digital diary that includes data visualization could combine the benefits of both annotation methods.
However, the study also showed that participants can find it difficult to interpret the acceleration data correctly and thus set inaccurate annotations. As our trained models show, this also has a strong influence on the classification result. If such a tool is to be made available to study participants, it must be ensured that they have the necessary knowledge and tools to be able to interpret these data. Thus, to ensure the success of future long-term and real-world activity recognition projects, prior training of the study participants regarding data interpretation is of crucial importance if a data visualization is supposed to be used.
Apart from trying to solve annotation difficulties during the annotation phase itself, we can also partially counter wrong or noisy classified data by using machine learning techniques like Bootstrapping, see Miu et al. [24] or using a loss function that specifically tries to counteract this problem, such as [5, 25]. By using Bootstrapping, the machine or deep learning classifier is initially trained by a small subset of high-confident labels and further improved by using additional data. However, this technique comes with the trade-off that whenever wrong-labeled data is introduced as training data, the error will get propagated into the model. An effect that sooner or later occurs as long as the annotation methodologies themselves are not further researched. Other machine learning techniques that can work with noisy labels, see [26], are already successfully tested for Computer Vision problems and can, in theory, be adopted for Human Activity Recognition. However, earlier research has shown that not every technique that is applicable in other fields is also applicable to sensor-based data [16].
5.1 Discussing different Annotation Biases
We need to take into account that several biases could have been introduced due to the chosen annotation guidelines and tools. For example, the usage of in situ annotation methods during the day can have a positive effect on the self-recall capabilities of a participant at the end of the day. However, the comparison of consistencies across methods does not confirm that this effect indeed occurred. Every study participant showed an almost complete overall profile of self-recall annotations, even though the person has not used or has incomplete in-situ annotations, see Figure 4. However, deeper investigations are needed to be able to understand such effects better.
[27] lists the following 3 biases for sensor-based human activity data: Self-Recall bias [28], Behavior bias [3] and the Self-Annotation bias [27]. We showed that indeed a time-deviation bias (which can be seen as a self-recall bias) has been introduced to annotations created with the pure self-recall method , and that such a bias affects the classifier negatively. However, visualizing the sensor data can counter this effect because it was easier for participants to detect active phases in hindsight.
A behavior bias can be neglected, because the participants were not monitored by a person or video camera during the day and the minimalistic setup of one wrist-worn smartwatch does not influence one’s behavior since the wearing comfort of such a device is generally perceived as positive [29]. A self-annotation bias, a bias that occurs if the annotator labels their data in an isolated environment and cannot refer to an expert to verify an annotation, did occur as well. With the deep learning analysis, we were able to show that the classifier was less negatively impacted by this bias than by time-deviation bias.
6 Conclusions
We argue that the annotation methodologies for benchmark datasets in Human Activity Recognition do not yet capture the attention it should. Data annotation is a laborious and time-consuming task that often cannot be performed accurately and conscientiously without the right tools. However, there is a very limited number of tools that can be used for this purpose and often they do not pass the prototype status.
Only a few scientific publications, such as [13], focus on annotations and their quality. However, the use of properly annotated data drastically affects the final capacities of the trained machine or deep learning model. Therefore, we consider our study to be important for the HAR community, as it analyzes this topic in greater depth and thus provides important insights that go beyond the current state of science. Table 4 summarizes the advantages and disadvantages of every method.
| Methodology | Advantages | Disadvantages | ||||||||||||
To guarantee high-quality annotations in future studies, especially in an uncontrolled real-world environment where a video recording is not available for the subsequent acquisition of ground truth, further research in this field is necessary. The methodologies used for annotating activity data need to be challenged and further developed.
The combination of a (handwritten) diary with a correction aided by a data visualization in hindsight shows the best results in terms of consistency and missing annotations and provides accurate start and end times. However, this combination results in additional work for the study participants and therefore, remains a trade-off between additional workload and annotation quality.
Lessons Learned. During this study, we gained insights about the effects of different annotation methods on the reliability and consistency of annotations and finally on the classifier itself, but also about training deep learning models on data recorded in-the-wild. In this chapter, we would like to share these insights to help other researchers perform their experiments more successfully. With regards to Table 4, we can narrow down specific study setups that either benefit more from self-recall or in situ annotation methods. As part of our annotation guidelines, we allowed our study participants to name their activities as they wished. Therefore, we were forced to simplify certain activities. To be able to create a real-world dataset that contains complex classes or even classes that consist of several subclasses, more elaborated annotation methods and tools must be developed. We believe that with the currently available resources, the hurdle lies very high for such datasets to be annotated accurately.
Our study includes people who cycle to work in their daily work routine and others who commute by public transport or work in a home office environment. Thus, each study participant has his or her set of daily repetitive activities. Due to the nature of our dataset as one recorded in a real-world and long-term scenario, the number of labeled samples is rather small, and given labels vary participant-dependent. This mix of factors creates a bias in the dataset and we concluded that a cross-participant train-/test-strategy is not appropriate for our study design and would not give meaningful insights, since every study participant has their own set of unique activities which are too different and hardly generalizable. Therefore, for certain studies, the commonly known and accepted Leave-One-Subject-Out Cross-Validation is not suitable.
Declarations
Conflict of Interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Funding
This project is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 425868829 and is part of Priority Program SPP2199 Scalable Interaction Paradigms for Pervasive Computing Environments.
Ethics
The studies involving human participants were reviewed and approved by the Ethics Committee of the University of Siegen (ethics vote #ER_12_2019). The participants provided their written informed consent to participate in this study and were briefed prior to the data capture about the study goals.
Availability of data and materials
The datasets generated during and/or analyzed during the current study are available at the Zenodo repository: https://doi.org/10.5281/zenodo.7654684.
Code availability
All code is made publicly available on our GitHub repository https://github.com/ahoelzemann/annotationMatters.
Consent to participate
Informed consent was obtained from all individual participants included in the study.
Consent for publication
The participants have consented to the submission of the manuscript to the journal.
Authors’ contributions
Alexander Hoelzemann has performed the implementation and implemented all studies and visualizations, Kristof Van Laerhoven has guided this work and assisted in the methodologies. Both authors have contributed substantially to the writing of this manuscript.